OpenStudy (cmtboy2):

William began to solve an equation. Which expression below shows the next step William needs to take? William’s Solution:

OpenStudy (cmtboy2):

A. 2.4 + 8 B. 2.4 – 8 C. 2.4 • 8 D. 2.4 ÷ 8

OpenStudy (thesmartestderpalive):

Whats Williams solution?

OpenStudy (cmtboy2):

8*m=2.4 8/ 8*m=?

OpenStudy (dustinrathke24):

Divide both sides by 8. So you get 8m/8=2.4/8.

OpenStudy (dustinrathke24):

Meaning the next step is 2.4/8.

OpenStudy (cmtboy2):

8/8 is 1

OpenStudy (dustinrathke24):

That's right, so 8m/8 equals m.

OpenStudy (dustinrathke24):

Or 1m.

OpenStudy (cmtboy2):

2.4/8=0.3

OpenStudy (dustinrathke24):

Yeah, but I think it just wants to know that step.

OpenStudy (cmtboy2):

so it wold be d

OpenStudy (dustinrathke24):

Yes.

OpenStudy (cmtboy2):

thank you

OpenStudy (dustinrathke24):

No problem.

OpenStudy (cmtboy2):

i got more

OpenStudy (dustinrathke24):

K.

OpenStudy (cmtboy2):

What is the value of d? 11d = –44 d =

OpenStudy (cmtboy2):

-144

OpenStudy (dustinrathke24):

You need to get d=something, so divide both sides by the coefficient 11.

OpenStudy (dustinrathke24):

11d/11=-44/11 d=-4

OpenStudy (cmtboy2):

11/11 is 1

OpenStudy (dustinrathke24):

Right, and 1d is the same thing as d. So that gives you d=-44/11

OpenStudy (cmtboy2):

so the answer is -44

OpenStudy (dustinrathke24):

No, -44 divided by 11. / means divide by

OpenStudy (cmtboy2):

oh so divid 11 by -44 that =-4

OpenStudy (dustinrathke24):

Yeah, but it's -44 divided by 11. You said it backwards.

OpenStudy (cmtboy2):

oh so now the answer is -4 right

OpenStudy (cmtboy2):

Which equations have a negative solution? Choose all answers that are correct.